OpenStudy (anonymous):

Admission to the movies costs $8 for adults and $4 for kids. The total bill for your entire extended family was $72. This scenario can be modeled by the equation 8a + 4k = 72. Which is not a possible combination of adults and kids? 3 adults and 12 kids 5 adults and 6 kids 7 adults and 4 kids 8 adults and 2 kids

OpenStudy (anonymous):

I know this is easy but just getting a 2nd opinion

OpenStudy (anonymous):

B

jimthompson5910 (jim_thompson5910):

Plug each option into the equation so for option A), we are given a = 3 and k = 12 8a + 4k = 72 8*3 + 4*12 = 72 24 + 48 = 72 72 = 72 so option A) is a solution. We want nonsolutions, so we eliminate choice A

OpenStudy (anonymous):

Yeah i knwo I did all of them and b equals 64 so it is not a solution

jimthompson5910 (jim_thompson5910):

good, you got it

OpenStudy (anonymous):

yeah thanks for all of ya'll's help :)

jimthompson5910 (jim_thompson5910):

you're welcome